OpenStudy (anonymous):

Which algebraic expression would not be equivalent if the number and the variable were reversed? n+8 m-5 kx8 4+x

OpenStudy (anonymous):

Test each one: \[n+8=8+n\] TRUE \[m-5=5-m\] FALSE \[k8=8k\] TRUE \[4+x=x+4\] TRUE Only B is not true when swapped.